In( )-Athz, , i = 1,2, ..,8 Age(z) Sample size 100 # of sickness | | 16 30 35 100 31 40 45 100 50 50 60 100 81 65 100 92 100 100 100 43 62 72 After the computation, we have β0--45 and β,-0.10
Answer #1

Since \alpha =0.05, therefore, 1-\alpha =0.95, you need to determine 95% confidence interval.

Now put the data in SPSS data editor window, then use the following command

Analyze

    Regression

       Binary Logistic

          Dependent: AGE

          Covariate: NOSICK

           OK.

Then you find the following output result

You have seen that the standard error of AGE, i.e. \beta1 is 0.008. Therefore the 95% confidence interval will be

\beta_{1} -Z_{0.025}se\left ( \beta _{1} \right )\leqslant \beta _{1} +Z_{0.025}se\left ( \beta _{1} \right )

=0.10-1.96\ast 0.008\leqslant \beta _{1}\leqslant 0.10+1.96\ast 0.008

=0.084\leqslant \beta _{1}\leqslant 0.12

Hence the required confidence interval is (0.084,0.12).